OpenStudy (anonymous):

Use Newton’s method to find an approximate root (accurate to six decimal places). Sketch the graph and explain how you determined your initial guess. cos x - x = 0

OpenStudy (anonymous):

@wio can u help?

OpenStudy (anonymous):

Okay, do you know how to do Newton's method?

OpenStudy (anonymous):

yea this rigth? x n + 1 = xn - f(xn)/f'(xn)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what is your derivative?

OpenStudy (anonymous):

okay i got (sin (x) + 1 = 0 )

OpenStudy (anonymous):

Close, but you don't need the = 0 part.

OpenStudy (anonymous):

where do you want to start?

OpenStudy (anonymous):

What will be your initial guess?

OpenStudy (anonymous):

so its sin(x) + 1?

OpenStudy (anonymous):

\[ (\cos x-x)'=-\sin x - 1 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what is f(x)/f'(x)?

OpenStudy (anonymous):

so the initial will be 1?

OpenStudy (anonymous):

Sure, it doesn't matter too much.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what comes after 1?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

lol oh you meant the next step?

OpenStudy (anonymous):

No, I mean, what is \[ 1+\frac{\cos x-x}{\sin x +1} \]

OpenStudy (anonymous):

where \(x=1\)

OpenStudy (anonymous):

ohh i got 0.750

OpenStudy (anonymous):

just keep going

OpenStudy (anonymous):

umm ok

OpenStudy (anonymous):

once the \[ f(x)/f'(x) \]term is under 6 decimals, you can stop.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so 0.750363

OpenStudy (anonymous):

need to keep going

OpenStudy (anonymous):

ok 0.7503638678

OpenStudy (anonymous):

0.750363867840243893034942306682176853246993065855359030966583... lol

OpenStudy (anonymous):

You're not getting closer to the actual answer for some reason

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cos+x+-+x=0

OpenStudy (anonymous):

this is the answer

OpenStudy (anonymous):

so the actual answer is 0

OpenStudy (anonymous):

\[x \approx 0.739085\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok